OpenStudy (anonymous):

A textbook search community is considering 17 books for possible adoption. The committee decided to send only 8 out of the 17...in how many ways can it do so

OpenStudy (dls):

C(17,8)

OpenStudy (anonymous):

17/8 17 -8 9 after these steps what do I do?

OpenStudy (dls):

\[\Huge C(17,8) = \frac{17!}{9! 8!}\]

OpenStudy (anonymous):

count down to 8 from 17?

OpenStudy (kropot72):

These are different groups of books, each group being in no particular order. So you want the number of combinations of the 17 books taken 8 at a time. \[17C8=\frac{17!}{8!9!}=\frac{17\times16\times15\times14\times13\times12\times11\times10}{8\times7\times6\times5\times4\times3\times2}\]

OpenStudy (dls):

17! = 17x16x15x14x13x12x11x10x9!